OpenStudy (anonymous):

Approximately how many fewer people live in Arica than in Europe and North America combined? Africa 971500000 Europe 689650000 North America 330850000 answer choices are: 48,750,000 33,800,000 97,175,000 33,085,000

OpenStudy (anonymous):

europe and north america combined=689650000+330850000=1020500000. so (E+NA)-Africa=1,020,500,000-971,500,000 =?

OpenStudy (anonymous):

This is simply done by adding the populations of Europe and North America, then subtracting the population of Africa from it.